OpenStudy (anonymous):

Solve -36 = j + 13 for j.

OpenStudy (skullpatrol):

This -36 = j + 13 is an equation with the variable on one side right?

OpenStudy (skullpatrol):

The equation is asking what number j do you have to add to 13 to get -36?

OpenStudy (skullpatrol):

If you added the number -13 to 13 you would get 0. If you added the number -26 to 13 you would get -13. But we want the number -? added to 13 to get -36... try -40 added to 13 gives -(|-40|-|13|)= -27. Try -50 added to 13 gives -(|-50|-|13|)=-37 <---close!!!

OpenStudy (anonymous):

Solve -36 = j + 13 for j. -36 = j + 13 Subtract 13 from both sides. -36 - 13 = j + 13 - 13 -49 = j

OpenStudy (anonymous):

J=-49